# FacLens: Transferable Probe for Foreseeing Non-Factuality in Large Language Models

### v3

@@ 4.2 Transferability of FacLens
Why Explore the Transferability of FacLens Across LLMs. When it comes to multiple LLMs, the process of training data construction becomes resource-intensive and time-consuming, because each LLM needs to conduct costly response generation for the training data construction (see Section 3.2). In this subsection, we discover the transferability of FacLens, which allows us to label training data on just one LLM and transfer the FacLens to support other LLMs. Appendix B illustrates the more efficient labeling process enabled by the transferability of FacLens.
Notably, the transferability of FacLens is inspired by insights from human cognition. Research has shown that individuals with similar cognitive styles and encoding strategies exhibit similar brain activity when performing specific tasks (Miller et al., 2012). In this analogy, we consider LLMs as individuals with similar cognitive styles and encoding strategies, given their common reliance on the Transformer architecture and overlapping training datasets. Accordingly, we assume that LLMs display similar cognitive patterns used by NFP.
Why Domain Adaptation is Effective for Transferring FacLens Across LLMs. Domain adaptation (DA) is an approach in transfer learning that transfers information from a source domain to improve performance in a target domain (Ben-David et al., 2006; Liu et al., 2022; Moreno-Torres et al., 2012). The premise of DA is that the source and target domains have distinct marginal probability distributions $P(\mathbf{X})$, but share similar conditional probability distributions $P(\mathbf{y}|\mathbf{X})$ (i.e., no significant concept shift) (Liu et al., 2022; Moreno-Torres et al., 2012). In this paper, we refer to the domain as,
@@
Figure 3: Performance comparison between the FacLens $f_{m}$ trained on an individual domain and the FacLens $f_{mix}$ trained on the mixture domain. Each LLM corresponds to an individual domain.
Figure 4: Distribution of prediction gap $\delta$ between individual-domain FacLens and mixture-domain FacLens over all questions (including questions from PQ, EQ, and NQ).
Figure 5: Visualization of NFP features extracted from hidden question representations across LLMs, where the blue and dark yellow points denote the positive and negative samples, respectively.
Different domains have different $P(\mathbf{X})$. If $P(\mathbf{y}|\mathbf{X})$ of different domains exhibit similar forms, we can say that similar NFP patterns exist in the hidden question representations sourced from different LLMs, and thus we can conduct unsupervised domain adaptation to derive a cross-LLM FacLens.
@@
$P_{mix}(\mathbf{X},\mathbf{y})=\sum\nolimits_{i=1}^{M}\alpha_{i}\cdot P_{m_{i}}(\mathbf{X},\mathbf{y})\quad s.t.\quad\sum\nolimits_{i=1}^{M}\alpha_{i}=1$ (2)
@@
where $m_{i}$ denotes the $i$-th LLM, $10<\alpha_{i}<$ 1 represents the proportion of $D_{i}$ in the mixture domain, and $M$ is the number of individual data domains (i.e. the number of LLMs). Here we set $\alpha_{i}=\frac{1}{M}$.
@@
Therefore, the FacLens trained on the mixture domain follows the following conditional distribution,
@@
$P_{mix}(\mathbf{y}|\mathbf{X})=\sum\nolimits_{i=1}^{M}\beta_{i}(\mathbf{X})\cdot P_{m_{i}}(\mathbf{y}|\mathbf{X}),\quad\beta_{i}(\mathbf{X})=\frac{\alpha_{i}\cdot P_{m_{i}}(\mathbf{X})}{\sum_{j=1}^{M}\alpha_{j}\cdot P_{m_{j}}(\mathbf{X})}$ (3)
@@
It is readily derived that $\sum\nolimits_{i=1}^{M}\beta_{i}(\mathbf{X})=1$, and $10<\beta_{i}(\mathbf{X})<$ 1 if $P_{m_{1}}(\mathbf{X}),P_{m_{2}}(\mathbf{X}),\cdots,P_{m_{M}}(\mathbf{X})$ are not disjoint. Therefore, if there is no concept shifts between individual data domains, we have,
$P_{mix}\left(\mathbf{y}|\mathbf{X}\right)=P_{m_{1}}\left(\mathbf{y}|\mathbf{X}\right)=P_{m_{2}}\left(\mathbf{y}|\mathbf{X}\right)=\cdots=P_{m_{M}}\left(\mathbf{y}|\mathbf{X}\right)$ (4)
@@
Conversely, if significant concept shifts exist between individual domains, Eq. 4 is not valid, as there must exists at least a domain $D_{i}$ where $P_{mix}\left(\mathbf{y}|\mathbf{X}\right)$ is very different from $P_{m_{i}}\left(\mathbf{y}|\mathbf{X}\right)$. Consequently, on the test set of domain $D_{i}$, $f_{mix}$ will noticeably underperform $f_{m_{i}}$, where $f_{mix}$ is trained on data of $D_{mix}$ and $f_{m_{i}}$ is trained on data of $D_{i}$. For simplicity, we use $f_{m}$ to denote a FacLens trained on an individual domain.
@@
Observation 1. As our experiments consider four popular LLMs, we have four individual domains. Each individual domain has its training, validation, and test sets. The training sets of all the individual domains form the training set of the mixture domain. Notably, the hidden dimension of Qwen2-1.5B-Instruct is different from that of the other three LLMs, so we introduce an additional linear layer to reshape the Qwen2’s hidden question representations to match the dimension of the other three LLMs. In Figure 3, $f_{mix}$ exhibits comparable performance to $f_{m}$ on the test set of each individual domain, indicating similar $P(\mathbf{y}|\mathbf{X})$ across individual domains, i.e., there is no significant concept shifts between individual domains.
@@
Observation 2. We further assess the degree of concept shift between an individual domain and the mixture domain by the distribution of prediction gap $\delta=||\mathbf{p}^{m}(y=1|\mathbf{x})-\mathbf{p}^{mix}(y=1|\mathbf{x})||$, where $\mathbf{p}^{m}$ is calculated by $f_{m}$, while $\mathbf{p}^{mix}$ is calculated by $f_{mix}$. Figure 4 shows that the values of $\delta$ are concentrated around zero, i.e., $P_{m_{1}}(\mathbf{\mathbf{y}|X}),P_{m_{2}}(\mathbf{\mathbf{y}|X}),\cdots,P_{m_{M}}(\mathbf{\mathbf{y}|X})$ and $P_{mix}(\mathbf{\mathbf{y}|X})$ are likely to have similar forms.
@@
Observation 3 (Visualization). To further demonstrate the feasibility of domain adaptation for FacLens, we visualize the NFP features from different domains. Given hidden question representations sourced from different LLMs, we use the encoder of $f_{mix}$ to extract the NFP features and visualize them with t-SNE. The visualization result is shown in Figure 5, where the blue points denote the positive samples while the dark yellow points denote the negative samples. Although these points are sourced from different LLMs, we can see that a unified classification boundary can be applied to them, thereby further demonstrating the similar $P(\mathbf{y}|\mathbf{X})$ across individual domains.
The above observations form the cornerstone of unsupervised domain adaptation (Zhuang et al., 2021; Ben-David et al., 2006; Kouw & Loog, 2021) for cross-LLM FacLens, highlighting the transferability of FacLens across different LLMs.
@@ 4.3 Cross-LLM FacLens
We have verified that $P_{m_{i}}(\mathbf{X})\neq P_{m_{j}}(\mathbf{X})$ and $P_{m_{i}}(\mathbf{y}|\mathbf{X})\approx P_{m_{j}}(\mathbf{y}|\mathbf{X})$, indicating that we can use domain adaptation (DA) to train a cross-LLM FacLens for LLM $m_{j}$ leveraging label information from LLM $m_{i}$. The distribution shift between a source domain $D_{S}$ and a target domain $D_{T}$, is due to the difference of LLMs. $D_{S}$ has labeled data, yet $D_{T}$ has no label information.
@@
Unsupervised Domain Adaptation for Cross-LLM FacLens. We adopt the classic Maximum Mean Discrepancy (MMD) loss (Gretton et al., 2012) to find a domain-invariant NFP feature space, based on which FacLens predicts the labels. The MMD loss calculates the distance between two distributions in the reproducing kernel Hilbert space (RKHS) (Smola et al., 2007). We denote the NFP features in the source and target domains as $Z_{S}=\{\mathbf{z}_{S,i}\}_{i=1}^{N_{S}}$ and $Z_{T}=\{\mathbf{z}_{T,j}\}_{j=1}^{N_{T}}$, respectively. The encoder $g_{enc}$ in FacLens is optimized by minimizing the MMD loss.
@@
$\mathcal{L}_{\text{MMD}}\left(Z_{S},Z_{T}\right)=\frac{1}{N_{S}^{2}}\sum_{i,j=1}^{N_{S}}k(\mathbf{z}_{S,i},\mathbf{z}_{S,j})+\frac{1}{N_{T}^{2}}\sum_{i,j=1}^{N_{T}}k(\mathbf{z}_{T,i},\mathbf{z}_{T,j})-\frac{2}{N_{S}N_{T}}\sum_{i=1}^{N_{S}}\sum_{j=1}^{N_{T}}k(\mathbf{z}_{S,i},\mathbf{z}_{T,j})$ (5)
@@
where $\mathbf{z}_{S,i}=g_{enc}\left(\mathbf{x}_{S,i}\right),\mathbf{z}_{T,j}=g_{enc}\left(\mathbf{x}_{T,j}\right)$, $N_{S}=N_{T}=|\mathcal{Q}_{train}|$ is the number of questions for training, and $k\left(\cdot\right)$ denotes a kernel function. We discuss the choice of kernel function in Appendix G. The hidden question representations are taken from the middle layer of the LLM.
@@
Importantly, we also use the CE loss to optimize $g_{enc}$ and $g_{clf}$ with the labeled data in $D_{S}$, which collaborates the MMD loss to find the latent feature space for NFP.
$\mathcal{L}_{\text{DA}}\left(D_{S},D_{T}\right)=\mathcal{L}_{\text{MMD}}\left(Z_{S},Z_{T}\right)+\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\mathcal{L}_{\text{CE}}\left(g_{clf}\left(\mathbf{z}_{S,i}\right),y_{S,i}\right)$ (6)
@@
Notably, if LLMs have distinct hidden dimensions, we introduce an additional linear layer to reshape the target domain’s hidden question representations to match the dimension of the source domain’s hidden question representations. We demonstrate that FacLens can transfer across LLMs of distinct hidden dimensions in Figure 6 and Appendix H.
Question-Aligned Strategy for Mini-Batch Training. MMD loss involves multiplications of feature matrices, which can lead to out-of-memory on GPUs when the number of training instances exceeds a certain threshold. To address this issue, mini-batch training is used by cross-LLM domain adaptation of FacLens. Specifically, in each mini-batch, two sets of questions $\mathcal{\overline{Q}}_{S}$ and $\mathcal{\overline{Q}}_{T}$ are sampled from $Q_{train}$ to empirically approximate the population distributions of the source and target domains. This raises a question: should $\mathcal{\overline{Q}}_{S}$ and $\mathcal{\overline{Q}}_{T}$ be identical?
Given a range of questions, the distribution $P(\mathbf{Z})$ is determined by the LLM. In a mini-batch, the number of sampled questions is limited, so the estimation of $P_{S}\left(\mathbf{Z}\right)$ and $P_{T}\left(\mathbf{Z}\right)$ within the mini-batch is likely to be affected by the sampling process. Hence, we propose to use the same set of questions in each mini-batch, i.e., $\mathcal{\overline{Q}}_{S}=\mathcal{\overline{Q}}_{T}$, to alleviate the influence of sampling process in estimating the true distance between $P_{S}\left(\mathbf{Z}\right)$ and $P_{T}\left(\mathbf{Z}\right)$. We name the strategy question-aligned mini-batch training.
5 Experiments
5.1 Experimental Setup
We compare FacLens with existing NFP methods, which have been introduced in Section 2, including Entity-Popularity (Mallen et al., 2023), SAT Probe (Yüksekgönül et al., 2024), Self-Familiarity (Luo et al., 2023), and Self-Evaluation (Kadavath et al., 2022). As Self-Evaluation fully fine-tunes the LLM for NFP, we adopt LoRA (Hu et al., 2022) as a baseline to conduct parameter-efficient fine-tuning for NFP. We additionally consider a Prompting-based method, which directly asks the LLM whether the LLM knows the factual answer to the given question. Moreover, inspired by using the perplexity to evaluate the factual precision of responses (Min et al., 2023a), we consider perplexity (PPL) on the input question as a baseline (see Appendix C). [^4] We extend the calculation of PPL to be conducted in each layer to obtain multiple PPL values for a text. We determine the layer based on the performance on labeled data. We find that PPL prefers the last few layers. PPL measures how well a language model (LM) predicts a given text. As a low PPL value suggests that an LM has likely learned relevant texts, we regard PPL on the input question as a non-factuality predictor. Due to space limitation, we provide the hyper-parameter settings in Appendix D. As the number of positive samples is larger than that of negative samples (see Table 3), we adopt AUC, a widely used metric for imbalanced binary classification, as the evaluation metric.
5.2 Experimental Results
Probing hidden question representations for NFP demonstrates promising performance. In Table 1, FacLens exhibits promising performance compared to the baselines. For example, FacLens (last token, middle layer) reaches 85%+ AUC on PQ across different LLMs. Self-Evaluation and LoRA, regarded as two special representation-based NFP methods, also derive good performance. FacLens trains with a small number of parameters, achieving results comparable to or even better than those of Self-Evaluation and LoRA. Compared with SAT Probe, FacLens demonstrates that, in addition to attention weights, hidden representation is also useful for the NFP task.
Question-level modeling is more effective than focusing on queried entities. Entity-Popularity and Self-Familiarity focus on analyzing the queried entities. SAT Probe focuses on the constraint tokens, which include the entity tokens. Besides, researchers have found that LLM’s response can be embedded in the representation of queried entities, while they have not evaluated the effectiveness of hidden entity representations for NFP. Following their idea, we introduce FacLens-ent, which feeds the hidden representation related to entity tokens to FacLens. We use the Stanza NLP Package (Qi et al., 2020) to identify entities like persons, locations, and organizations in a question. In Table 1, FacLens consistently outperforms FacLens-ent, Entity-Popularity, Self-Familiarity, and SAT Probe, implying that focusing on the entities while overlooking the comprehension of the entire question could misguide the predictions.
LLMs have (mostly) known whether they know in their middle layers. An intriguing finding in Table 1 is that hidden question representations in the middle layer tend to be more beneficial compared to that in the last few layers. For this phenomenon, we conjecture that an LLM could already known “whether it knows” in its middle layer. As LLMs progress from the middle layer to the final layer, they may focus more on how to better organize their knowledge for final response. It is noteworthy that FacLens favors the last few layers of Qwen2. This may be attributed to the smaller scale of Qwen2-1.5B compared to the other three LLMs, which might necessitate additional layers to ascertain “whether it knows”.
FacLens stands out for its efficiency. Taking the LLaMA2-PQ NFP dataset as an example, which comprises 2,272 questions for training, 1,136 questions for validation, and 7,952 questions for testing. Table 2 reports the averaged training time per epoch and averaged prediction time per question of each method. In Table 1, Self-Evaluation and LoRA shows competitive performance. However, we can see that fine-tuning an LLM incurs significant computational costs. Importantly, if the LLM has been fine-tuned for a specific task, its ability on other tasks can be compromised (Yang et al., 2024b). As for the prediction, FacLens runs much faster than Self-Familiarity because Self-Familiarity involves multiple conversations with the LLM. Because FacLens utilizes the middle-layer hidden question representations, it runs faster than Prompting, PPL, LoRA, and Self-Evaluation which involve more layers in the LLM. SAT Probe extracts attention weights across all layers and attention heads, so FacLens runs faster than SAT Probe during prediction.
Table 1: Prediction performance of different NFP methods (AUC %).
@@ 5.2 Experimental Results
  • “–” means the method is not suitable for the QA dataset. We give the detailed explanation in the appendix E. “avg” refers to the averaged hidden representation of the input entities’ tokens or a question’s tokens. “Last token” refers to the hidden representation of the final token in the input entities or the question. The question consists of a chat template and the original question, where the chat template can prompt the LLM to better respond. Due to space limitation, we show the performance of FacLens (avg) in Appendix F.
Table 2: Efficiency evaluation of FacLens (seconds).
    Training-Free     Transferable     Training Time Per Epoch (Avg. )     Prediction Time Per Question (Avg. )

### v4

@@ 4.2 Transferability of FacLens
When it comes to multiple LLMs, the construction of NFP training data becomes resource-intensive and time-consuming, because each LLM needs to conduct costly response generation (see Section 3.2). Fortunately, we discover the transferability of FacLens, which allows us to label training data for just one LLM and adapt the FacLens to support other LLMs. In Appendix B, we illustrate the more efficient process of training data construction enabled by the transferability of FacLens.
Why Domain Adaptation is Effective for Transferring FacLens Across LLMs. Domain adaptation (DA) is an approach in transfer learning that transfers information from a source domain to improve performance in a target domain (Ben-David et al., 2006; Liu et al., 2022; Moreno-Torres et al., 2012). The premise of DA is that the source and target domains have distinct marginal probability distributions $P(\mathbf{X})$, but share similar conditional probability distributions $P(\mathbf{y}|\mathbf{X})$ (i.e., no significant concept shift) (Liu et al., 2022; Moreno-Torres et al., 2012). Here we refer to the domain as,
@@
Different domains naturally have different $P(\mathbf{X})$. If $P(\mathbf{y}|\mathbf{X})$ of different domains exhibit similar forms, we can perform DA to apply FacLens to other LLMs without new labels for training.
@@
$\begin{split}P_{mix}(\mathbf{X},\mathbf{y})&=\sum\nolimits_{i=1}^{M}\alpha_{i}\cdot P_{m_{i}}(\mathbf{X},\mathbf{y})\\ &\quad\text{s.t.}\quad\sum\nolimits_{i=1}^{M}\alpha_{i}=1\end{split}$ (2)
@@
where $M$ is the number of individual data domains (i.e. the number of different LLMs), $m_{i}$ denotes the $i$-th LLM, and $0<\alpha_{i}<1$ represents the proportion of $D_{i}$ in the mixture domain. Here we set $\alpha_{i}=\frac{1}{M}$.
@@
Therefore, the FacLens trained on the mixture domain follows the conditional distribution,
@@
$\begin{split}P_{mix}(\mathbf{y}|\mathbf{X})&=\sum\nolimits_{i=1}^{M}\beta_{i}(\mathbf{X})\cdot P_{m_{i}}(\mathbf{y}|\mathbf{X}),\\ \beta_{i}(\mathbf{X})&=\frac{\alpha_{i}\cdot P_{m_{i}}(\mathbf{X})}{\sum_{j=1}^{M}\alpha_{j}\cdot P_{m_{j}}(\mathbf{X})}\end{split}$ (3)
@@
It is readily derived that $\sum\nolimits_{i=1}^{M}\beta_{i}(\mathbf{X})=1$, and $0<\beta_{i}(\mathbf{X})<1$ if $P_{m_{1}}(\mathbf{X}),P_{m_{2}}(\mathbf{X}),\cdots,P_{m_{M}}(\mathbf{X})$ are not disjoint. If there are no concept shifts between individual data domains, we have,
$\begin{split}P_{mix}\left(\mathbf{y}|\mathbf{X}\right)&=P_{m_{1}}\left(\mathbf{y}|\mathbf{X}\right)=P_{m_{2}}\left(\mathbf{y}|\mathbf{X}\right)\\ &=\cdots=P_{m_{M}}\left(\mathbf{y}|\mathbf{X}\right)\end{split}$ (4)
@@
Conversely, if significant concept shifts exist between individual domains, Eq. 4 is not valid, as there must exist at least a domain $D_{i}$ where $P_{mix}\left(\mathbf{y}|\mathbf{X}\right)$ is very different from $P_{m_{i}}\left(\mathbf{y}|\mathbf{X}\right)$. Consequently, on the test set of domain $D_{i}$, $f_{mix}$ will noticeably underperform $f_{m_{i}}$, where $f_{mix}$ is trained in $D_{mix}$, and $f_{m_{i}}$ is trained in $D_{i}$. For simplicity, we use $f_{m}$ to denote a FacLens trained on an individual domain.
Figure 3: Performance comparison between $f_{m}$ and $f_{mix}$. Similar performance suggests no significant concept shift across different domains. Trends are consistent on EQ and NQ (See Appendix D.2)
Figure 4: Distribution of $\delta$ between an individual-domain FacLens and the mixture-domain FacLens overall questions (questions from PQ, EQ, and NQ).
Figure 5: Visualization of NFP features extracted from different LLMs’ hidden question representations, where each domain corresponds to an LLM, and the blue and dark yellow points denote the positive and negative samples, respectively. Trends are consistent on EQ and NQ (see Appendix D.2).
@@
Given a specific fact-seeking QA dataset, we have four individual domains as our experiments consider four LLMs. Each individual domain has its training, validation, and test sets. The training sets of all the individual domains form the training set of the mixture domain. Notably, the hidden dimension of Qwen2-1.5B-Instruct is different from that of the other three LLMs, so we introduce an additional linear layer to reshape Qwen2’s hidden question representations to match the dimension of the other three LLMs.
Observation 1. Figure 3 shows the results on the PQ dataset, and we can observe that $f_{mix}$ exhibits comparable performance to $f_{m}$ on the test set of the corresponding individual domain, indicating similar $P(\mathbf{y}|\mathbf{X})$ across different domains. Hence, we can conduct DA to derive a FacLens used by different LLMs.
@@
Observation 2. We measure concept shift between an individual domain and the mixture domain by $\delta=||\mathbf{p}^{m}(y=1|\mathbf{x})-\mathbf{p}^{mix}(y=1|\mathbf{x})||$, where $\mathbf{p}^{m}$ is computed by $f_{m}$ and $\mathbf{p}^{mix}$ by $f_{mix}$. The values of $\delta$ are mostly near zero (see Figure 4), indicating that $P_{m_{1}}(\mathbf{y|X}),P_{m_{2}}(\mathbf{y|X}),\dots,P_{m_{M}}(\mathbf{y|X})$ and $P_{mix}(\mathbf{y|X})$ have similar distributions.
@@
Observation 3. Given hidden question representations from different LLMs, we use the encoder of $f_{mix}$ to extract the NFP features and visualize them with t-SNE. The results of PQ are shown in Figure 5, where the positive and negative samples are represented by points of different colors. Although these points are sourced from different LLMs, we can see that a unified classification boundary can be applied to them, further demonstrating the similar $P(\mathbf{y}|\mathbf{X})$ across different domains.
@@ 4.3 Cross-LLM FacLens
We have verified that we can use DA to train a cross-LLM FacLens for LLM $m_{j}$ leveraging label information from LLM $m_{i}$. Here the distribution shift between a source domain $D_{S}$ and a target domain $D_{T}$, is due to the difference of LLMs. $D_{S}$ has labeled data, yet $D_{T}$ has no label information.
@@
Unsupervised Domain Adaptation. The classic Maximum Mean Discrepancy (MMD) loss (Gretton et al., 2012) is used to find a domain-invariant NFP feature space, based on which FacLens predicts the labels. The MMD loss calculates the distance between two distributions in the reproducing kernel Hilbert space (RKHS) (Smola et al., 2007). We denote the NFP features in the source and target domains as $Z_{S}=\{\mathbf{z}_{S,i}\}_{i=1}^{N_{S}}$ and $Z_{T}=\{\mathbf{z}_{T,j}\}_{j=1}^{N_{T}}$, respectively, where $\mathbf{z}_{S,i}=g_{enc}\left(\mathbf{x}_{S,i}\right)$, and $\mathbf{z}_{T,j}=g_{enc}\left(\mathbf{x}_{T,j}\right)$. The encoder $g_{enc}$ in FacLens is optimized by minimizing the MMD loss,
@@
$\begin{split}\mathcal{L}_{\text{MMD}}\left(Z_{S},Z_{T}\right)&=\frac{1}{N_{S}^{2}}\sum_{i,j=1}^{N_{S}}k(\mathbf{z}_{S,i},\mathbf{z}_{S,j})\\ &\quad+\frac{1}{N_{T}^{2}}\sum_{i,j=1}^{N_{T}}k(\mathbf{z}_{T,i},\mathbf{z}_{T,j})\\ &\quad-\frac{2}{N_{S}N_{T}}\sum_{i=1}^{N_{S}}\sum_{j=1}^{N_{T}}k(\mathbf{z}_{S,i},\mathbf{z}_{T,j})\end{split}$ (5)
@@
where $N_{S}=N_{T}=|\mathcal{Q}_{train}|$ is the number of questions for training, and $k\left(\cdot\right)$ denotes a kernel function. We extract hidden question representations from the LLM’s middle layer.
@@
Importantly, we also use the CE loss to optimize $g_{enc}$ and $g_{clf}$ with the labeled data in $D_{S}$, which collaborates the MMD loss to find the latent feature space for NFP. Finally, the loss function is,
$\mathcal{L}_{\text{DA}}=\mathcal{L}_{\text{MMD}}+\frac{1}{N_{S}}\sum_{i=1}^{N_{S}}\mathcal{L}_{\text{CE}}\left(g_{clf}\left(\mathbf{z}_{S,i}\right),y_{S,i}\right)$ (6)
@@
Notably, if LLMs have distinct hidden dimensions, we introduce an additional linear layer to reshape the target domain’s hidden question representations to match the dimension of the source domain’s hidden question representations. We demonstrate that FacLens can transfer across LLMs of distinct hidden dimensions in Figure 6 and Appendix D.5. Besides, we discuss the choice of kernel function for MMD loss in Appendix D.6.
@@
  • “–” means the method is not suitable for the QA dataset. We give the detailed explanation in the appendix C.3. “avg” refers to the averaged hidden representation of the input entities’ tokens or a question’s tokens. “last token” refers to the hidden representation of the last token in the input entities or the question. The question consists of a chat template and the original question, where the chat template can prompt the LLM to better respond. Due to space limitation, we show the performance of FacLens (avg) in Appendix D.3.
Table 1: Prediction performance of different NFP methods (AUC %).
Question-Aligned Mini-Batch Training. In order to address GPU out-of-memory issues, FacLens employs mini-batch training for DA. In each mini-batch, we sample two question sets, $\mathcal{\overline{Q}}_{S}$ and $\mathcal{\overline{Q}}_{T}$, from $Q_{train}$, for two domains. This raises a question: are $\mathcal{\overline{Q}}_{S}$ and $\mathcal{\overline{Q}}_{T}$ identical? Given a range of questions, the distribution $P(\mathbf{Z})$ should be determined by the LLM. In a mini-batch, the number of sampled questions is limited, so the estimation of $P_{S}\left(\mathbf{Z}\right)$ and $P_{T}\left(\mathbf{Z}\right)$ within the mini-batch is likely to be affected by the sampling process. Hence, we propose to use the same question set for two domains in each mini-batch, i.e., $\mathcal{\overline{Q}}_{S}=\mathcal{\overline{Q}}_{T}$, to alleviate the influence of sampling process in estimating the true distance between $P_{S}\left(\mathbf{Z}\right)$ and $P_{T}\left(\mathbf{Z}\right)$.
5 Experiments
5.1 Experimental Setup
We compare FacLens with existing NFP methods, which have been introduced in Section 2, including Entity-Popularity (Mallen et al., 2023), SAT Probe (Yüksekgönül et al., 2024), Self-Familiarity (Luo et al., 2023), and Self-Evaluation (Kadavath et al., 2022). As Self-Evaluation fully fine-tunes the LLM for NFP, we adopt LoRA (Hu et al., 2022) as an additional baseline to conduct parameter-efficient fine-tuning. We also consider a Prompting-based method, which directly asks the LLM whether the LLM knows the factual answer to the given question. Moreover, inspired by using the perplexity to evaluate the factual precision of responses (Min et al., 2023b), we consider perplexity (PPL) on the input question as a baseline (see Appendix C.1). Due to space limitation, we provide the hyper-parameter settings in Appendix C.2. As the number of positive samples is larger than that of negative samples (see Table 3), we adopt AUC, a common metric for imbalanced binary classification, as the evaluation metric.
    Training-Free     Transferable
Training Time
Per Epoch (avg)
Prediction Time
Per Question (avg)